[The following information applies to the questions displayed below. Jeremy earned $100,000 in salary and $6,000 in interest income during the year. Jeremys employer withheld $11,200 of federal income taxes from Jeremys paychecks during the year. Jeremy has one qualifying dependent child who lives with him. Jeremy qualifies to file as head of household and has $23,000 in itemized deductions. (Use the tax rate schedules.) etermine Jeremys tax refund or taxes due Answer is complete but not entirely correct Amount $ 106,000 Description Gross income For AGI deductions Adjusted gross income ) Standard deduction ) Itemized deductions Greater of standard deduction or itemized deductions ) Taxable income $ 106,000 18,000 23,000 23,000 $ 83,000 $ 14,243 2,000 11,200 $ 4,298 ) Income tax liability Child tax credit ) Tax withholding Taxes refund[The following information applies to the questions displayed below. Jeremy earned $100,000 in salary and $6,000 in interest income during the year. Jeremys employer withheld $11,200 of federal income taxes from Jeremys paychecks during the year. Jeremy has one qualifying dependent child who lives with him. Jeremy qualifies to file as head of household and has $23,000 in itemized deductions. (Use the tax rate schedules.) Assume the original facts except that Jeremy had only $7.000 in itemized deductions. What is Jeremys tax refund or tax due? Description Amount (1) Gross income (2) For AGI deductions (3) Adjusted gross income (4) Standard deduction (5) Itemized deductions (6) Greater of standard deduction or itemized deductions (7) Taxable income (8) Income tax liability (9) Child tax credit (10) Tax withholding $ 106,000 $ 106,000 18,000 $ 106,000 Taxes due
Answer #1

Let's assume Jeremy to be an unmarried individual for the year 2018

Taxable income is $83000

Tax liability is calculated as follows:

up to 9525$ @10%=$952.5

above $9525 up to $38700 i.e., 38700-9525=29175 @12%=3501

above $38700 up to $ 82500 i.e., 82500-38700=43800@ 22%=9636

above $82500 up to $83000 i.e., 83000-82500=500@24%=120

total tax liability (952.5+3501+9636+120)=14209.5

gross income 106000

for AGI deductions

0
Adjusted gross total income 106000
standard deduction 18000
itemized deduction 23000
greater of the standardized and itemized deduction 23000
taxable income 83000
income tax liability 14210
child tax credit 2000
tax withholding 11200

taxes DUE

1010

IN THE SECOND CASE WHERE JEREMY HAD ONLY $7000 ITEMIZED DEDUCTION:

gross income 106000
for AGI deductions 0
Adjusted gross total income 106000
standard deduction 18000
itemized deduction 7000
greater of THE standardized and itemized deduction 18000
taxable income 88000
income tax liability 15410
child tax credit 2000
tax withholding 11200

taxes DUE

2210

NOTE: Tax rates change with assesses marital status, citizenship, and residential status and thus changes tax liability.

since the above are not mentioned in the given problem, they are assumed.